tareska:

The label on the car's antifreeze container claims to protect the car between −30°C and 130°C. To convert Celsius temperature to Fahrenheit temperature, the formula is C = 5 over 9 (F − 32). Write and solve the inequality to determine the Fahrenheit temperature range at which this antifreeze protects the car.

darkknight:

u got this @supie or me

supie:

ight so first the equation is C=5/9 (F-32 to find the Celcius when we got F so we gotta change it to F=C(9/5)+32 and then subtract both Celc. F=-30 (9/5)+32 =-22 F=130 (9/5)+32 =266 so it would be between -22°F and 266°F i think

darkknight:

yes, u are correct, @supie so yeah when you write your inequality it will be −22°F≤x≤266°F Where x is a Fahrenheit temperature at which this antifreeze protects the car.